A spinner is divde into 6 segments, each labled with a different positive integer. The theoretical probabilityof spinning a number less than 8 is 1. The theoretical probability of spinning a multiple of 3 is 1/3, a multiple of 4 is 1/6, and an odd number is 2/3. give me the numbers on the spinner

Answer by stanbon(75887)

A spinner is divde into 6 segments, each labled with a different positive integer. The theoretical probabilityof spinning a number less than 8 is 1. The theoretical probability of spinning a multiple of 3 is 1/3, a multiple of 4 is 1/6, and an odd number is 2/3. give me the numbers on the spinner
------------------------------------------------------------------------
P(Multiple of 3) = 1/3 = 2/6
3,6 are on the spinner
-----------------------------
P(Multiplo of 4) = 1/6
4 is on the spinner
-------------------------
P(odd) = 2/3 = 4/6
1,1,5 are on the spinner or 1,5,5 are on the spinner
------------------------------
The numbers on the spinner are: 1,1,3,4,5,6 or 1,3,4,5,5,6
--------------------------------------------------------------
Cheers
Stan H.